# Instability of a fluctuating biomimetic membrane driven by an applied uniform DC electric field

### v1

@@ IV.1 Paremeters effects
Next, we will illustrate the physical meaning at two limits $\beta\rightarrow 0,+\infty$. As $\beta\rightarrow 0$, there is no charge accumulated in the Debye layer, i.e., the Debye layer vanishes. The base state potential are constant in space as $\phi^{(0)}_{i}(z)=\pm V$, and the electric field $E_{i,z}^{(0)}(z)=0$. Thus, at this limit, the EK system behave like a LDM: without ion current ($g_{m}=0$), the membrane will be stable. Increasing $\beta$ represents increasing the effect of the membrane capacitance $C_{m}$, i.e. the ability to store charge at unite potential ($Q_{m}/V_{m}$), with $Q_{m}=|Q_{i}|$ (according to Eq. 6) as the macroscopic charge along the membrane on one side. As $\beta\rightarrow+\infty$, $V_{m}=0$, and the $Q_{m}$ also reaches its maximum as $\gamma\rho_{m}$, with $\rho_{m}|_{\beta\rightarrow+\infty}=2V/(1+\gamma^{2})$. The membrane capacitance $C_{m}=\varepsilon_{m}/d_{0}\rightarrow+\infty$ can be achieved by consider a membrane with zero-thickness $d_{0}\rightarrow 0$ but with finite $\varepsilon_{m}$. From an electrical standpoint, the membrane acts as a simple fluid–fluid interface: without potential jump across the membrane, the Debye layers on both sides still stores charges, and the internal normal traction from internal membrane disappear $\rho_{m}^{2}\gamma^{2}/\beta|_{\beta\rightarrow 0}=0$. In this limit, the Robin BCs in Eq. (15) reduces to the continuity of potential ${[\![}\phi{]\!]}=0$, and the continuity of the displacement field ${[\![}\varepsilon_{f}\bm{n\cdot E}{]\!]}=0$ at the interface. These two boundary conditions was used in an EK model for the simple fluid interface [22] that is impermeable to ions. It differs from the interfacial condition using ion conservation [32, 25]. Mechanically, the membrane is different from a simple interface, due to the surface-incompressibility resulting from the conservation of lipids. The limit at $g_{m}\rightarrow 0$ is also discussed in [10], which represents a short-circuited capacitor and the membrane acts as a simple fluid–fluid interface.
@@
In [10], the growth rate monotonically increases with $g_{m}$ and then saturates, while in current system, the growth rate first increases then decreases with $\beta$ and saturates to constant. This is due to the combination effect of internal and external effects. When excluding the internal normal traction (blue curve in Fig. 4(a)), the growth rate behaves monotonically. In Fig. 4(b), the external contribution to the growth rate increases with $\beta$ and saturates at a constant value, whereas the internal traction contribution initially rises from zero at $\beta=0$, peaks, and then decreases back to zero as $\beta\to\infty$. Specifically, $\lim_{\beta\to 0}s_{\text{in}}=\lim_{\beta\to\infty}s_{\text{in}}=0$, corresponding to the overlap between the full growth rate (black curve) and the growth rate excluding internal traction (blue curve) at these limits in Fig. 4(a). The internal traction contribution reaches its maximum at $\beta=\frac{\gamma}{1+\gamma}$, with $\max s_{\text{in}}=\frac{V^{2}\gamma^{3}}{1+\gamma}$. This peak results in the nonmonotonic behavior of the full growth rate in Fig. 4(a), which is further highlighted by the variation of the maximum growth rate shown in Fig. 4(c).
@@
Figure 5: (a) The effect of $\gamma$ on growth rate $s$ as a function of perturbation wavenumber $k$; (b) maximal growth rate $s_{max}$ as function of $\gamma$. We consider the system with $V=2$, $\beta=10$, and $Ca=2$.
@@
Fig. 5(a) shows the effect of the asymmetric conductivity $\gamma$ on the growth rate. Due to the reversability of the system, i.e. $s(\gamma)=s(1/\gamma)$, we only show results for $\gamma>1$. It is very anti-intuition that the membrane subjected to the fluids with the symmetric conductivity ($\gamma=1$) undergoes the most intense instability. As $\gamma$ increase from $1$, the maximum growth rate decrease as shown in Fig. 5(b), and the window for the unstable wavenumbers get narrowed. As $\gamma\rightarrow+\infty$, the system become stable. This result is very different from the known literature [10], where the instability increase with the asymmetry of conductivity. To understand the cause of this difference, it is essential to first examine the mechanism of the instability and how it diverges from the LDM framework. This will be explored in detail in the remainder of this section.
@@ IV.2.1 Stabilizing electrostatic body forces
The Coulomb body forces in the Stokes equation (18), $\bm{f}=-\rho\nabla\phi$, tend to move the positive charges $\rho$ in the direction of the electric field $\bm{E}$. In the base state, these body forces are balanced by the pressure gradient on both sides of the fluid, as evidenced by Eq. (26). Additionally, the flat interface remains free from normal traction arising from either electrostatic stress or elasticity, as well as surface tension.
Figure 6: $z$-component electric field $E_{z}$ (a,d) and charge density $\rho$ (b,e) for the base state and first order are used to illustrate the the first order body force component $\bm{f}^{(1)}\cdot\hat{e}_{z}=\rho^{(0)}E_{z}^{(1)}+\rho^{(1)}E_{z}^{(0)}$ in (c,f). We consider symmetric conductivity $\gamma=1$, $\beta=1$, $V=1$.
To understand the stabilizing effect indicated in Eq. (37) and Eq. (41), we illustrate the first order body force $\bm{f}^{(1)}=-(\rho^{(0)}\nabla\phi^{(1)}+\rho^{(1)}\nabla\phi^{(0)})$ in Fig. 6. The body force is calculated with a uni-modal perturbation with wavenumber $k=0.2$, corresponding to a wavelength 5 times the Debye length, i.e., $x=5$. The perturbation amplitude is set to $h=0.5$. It is expected that the body force acts only on the fluid in the Debye layer, decaying exponentially in the bulk. Within the Debye layer, the body force points downward at the interface’s crest and upward at the trough. This implies the restorative effect of the body force, consistent with the negative sign in the term Eq. (41) for small $k$.
@@ IV.2.2 Destabilizing external electrostatic surface traction
The external normal traction has the expression
${[\![}\partial_{z}\phi^{(0)}\partial_{z}\phi^{(1)}{]\!]}_{z=0}=(\partial_{z}\phi_{1}^{(0)}\partial_{z}\phi_{1}^{(1)}-\partial_{z}\phi_{2}^{(0)}\partial_{z}\phi_{2}^{(1)})|_{z=0}.$ (45)
In LDM framwork as in [10], when the membrane is fully charged, and no current cross the membrane, resulting in zero Omic current in the bulk fluid, the normal traction will be zero due to the absence of a base state electric field $\partial_{z}\phi^{(0)}_{i}=-E_{z}^{(0)}=0$. In the current study, the base state electric field is nonzero at the membrane due to the Debye layer $\partial_{z}\phi_{1}^{(0)}|_{z=0}=\partial_{z}\phi_{2}^{(0)}|_{z=0}=\gamma\rho_{m}$. Thus, Eq. (45) can be further simplified as
@@
[…]
In addition, we observe that in contrast to the LDM, the presence of a non-zero base state electric field $\bm{E}^{(0)}$ primarily contributes to the instability. In the LDM, a non-zero base state electric field necessitates either bulk current [10, 7] or membrane charging. However, within the EK framework, $\bm{E}^{(0)}$ consistently exists independently of other effects due to the Debye layer.
Figure 7: The uniform charge density at the base state (a) redistributes due to the perturbation $h$. (b) shows the perturbed charge density at first order and (c) shows the total charge denstity at the membrane as the summation of plot (a) and (b). The figure shows the example with $\gamma=1$, $\beta=1$, $V=1$.
Intuitively, we can understand such instability through the charge redistribution. Fig. 7 shows the charge distribution in the fluid along the membrane. At the base state (Fig. 7(a)), the charge is uniformly distributed along the membrane. However, when subjected to a small perturbation (Fig. 7(b)), the linear correction causes positive charge accumulation at the toughs on both sides of membrane, while negative charge accumulates at the crests. As a result (Fig.7(c)), on the upper lipid, positive charge accumulates denser at the troughs compared to the crests, while on the lower lipid, negative charge accumulates less at the troughs than at the crests. This accumulation will further result in a force which tends to enhance the perturbation. Similar mechanics of the instability resulting from surface charge redistribution due to the membrane undulation are also discussed in [15]. Different from the present study, the surface charge in  [15] are adsorbed on the interface, and such that compensates the charge distributed in Debye layer, i.e. each lipid layer is electroneutral.
Next, we aim to understand why the instability decreases with asymmetry of the conductivity. We consider the scenario where $\gamma\rightarrow 0$, i.e. $R=\frac{n_{1}^{\infty}}{n_{2}^{\infty}}\rightarrow\infty$. In this limit, $\rho^{(0)}_{1}|_{z=0}=\rho_{m}=\frac{2\beta V}{\gamma+(1+\gamma^{2})\beta}\rightarrow 2V$, $\rho^{(0)}_{2}|_{z=0}=-\gamma^{2}\rho_{m}\rightarrow 0$, and the net charge density across membrane ${[\![}\rho^{(0)}{]\!]}_{z=0}\rightarrow 2V$. However, the base state field $\partial_{z}\phi_{i}^{(0)}|_{z=0}=\gamma\rho_{m}\rightarrow 0$, resulting in the entire external traction term approaching zero. Therefore, the instability decreases as $\gamma\rightarrow 0$. Similar results can be derived as $\gamma\rightarrow\infty$, where ${[\![}\rho^{(0)}{]\!]}_{z=0}\rightarrow 2V$ and the base state field $\partial_{z}\phi_{i}|_{z=0}=\gamma\rho_{m}\rightarrow 0$. The result of these two limits is illustrated in Fig. 8, where the destabilizing external surface traction attains its maximum when subjected to the symmetric conductivity.
@@
Figure 8: The effect of the conductivity ratio $\gamma$ on (a) base state electric field $E_{z}^{(0)}|_{z=0}$, (b) the base state charage density accross the membrane, and (c) the external normal surface traction calculated according to Eq. (46). The example is given with $\beta=10$, $V=1$.
@@ IV.2.3 Destabilizing internal electrostatic surface traction
The electric stress induced by the rearranged charge at the membrane has also been used to explain the internal normal traction in [7], which establishes the system in the LDM framework, and the surface charge is brought to interface by the bulk current. In their work, the compressive electrostatic stress has the form $\varepsilon_{m}E_{m}^{2}/2$. For a perfectly flat membrane, the electric stress in the membrane is symmetrically balanced on both sides of the membrane. When subjected to a perturbation, the redistributed charge $\Sigma_{\pm}\sim\pm\varepsilon_{m}E_{m}$ results into an unbalanced stress $(\Sigma_{+}^{2}-\Sigma_{-}^{2})/\varepsilon_{m}$. This analysis shares the same situation as ours, where the the base state macroscopic charge is $Q_{i}^{(0)}=\pm E_{z}^{(0)}|_{z=0}=\pm\beta dE_{m}^{(0)}$.
@@
where the jump of the first order internal field $\bigg{[}\!\!\bigg{[}\partial_{z}\phi_{m}^{(1)}\bigg{]}\!\!\bigg{]}_{z=\pm d/2}\sim E_{m}^{(0)}dk^{2}h$ is nonzero due to the perturbation (see Appendix A for expression). Intuitively, we can understand this jump as the virtual charges inside the membrane, and driven by the field $E_{m}^{(0)}$. Like the external traction, we consider the deformation in physical domain, which has the same formulation but with $h=h(x)$. At the membrane where $h(x)<0$, the virtual charges are positive and driven by a negative internal field. Thus, the internal surface traction will point downward and vice versa, which renders the membrane unstable to deformations.
As discussed around Eq. (37), the internal normal traction in this work is in agreement with the total traction in [7], where the external traction vanishes in the limit of large $\varepsilon_{m}$ and small membrane conductivity $\sigma_{m}$. In their work, the internal field $E_{m}$ is related with the external field $E$ though the conservation of current, and thus $E_{m}\sim E\sigma_{i}\sigma_{m}$, and thus small $\sigma_{m}$ indicates a large internal field. In our work, the internal field $E_{m}^{(0)}\sim E_{z}^{(0)}|_{z=0}/(\beta d)$, and thus $\beta d\ll 1$ indicates a large fields ratio $E_{m}^{(0)}/E_{z}|_{z=0}$. Keeping $d$ finite and letting $\beta\rightarrow 0$, we see the destabilizing effect of the normal traction documented by the internal traction in Eq. (43). This conclusion is in agreement with [19, 29], where the “inside” contribution is in general dominant when large voltage drop occurs across the membrane, and in our work $\beta\rightarrow 0$ yields the largest transmembrane potential $V_{m}$.
Similar to the external surface traction, as $\gamma\rightarrow 0$ or $\gamma\rightarrow+\infty$, the base state external field $E_{z}^{(0)}|_{z=0}\rightarrow 0$ and thus $E_{m}^{(0)}\rightarrow 0$, and the interal traction term vanishes.

### v2

@@ IV.1 Effect of physical parameters
Next, we illustrate the physical meaning at two limits $\beta\rightarrow 0,+\infty$. As $\beta\rightarrow 0$, there is no charge accumulated in the Debye layer, i.e., the Debye layer vanishes. The base state potential are constant in space as $\phi^{(0)}_{i}(z)=\pm V$, and the electric field $E_{i,z}^{(0)}(z)=0$. Thus, at this limit, the EK system behaves like the LDM one: without ion current ($g_{m}=0$), the membrane is stable. Increasing $\beta$ represents increasing the effect of the membrane capacitance $C_{m}$, i.e. the ability to store charge at unit potential ($Q_{m}/V_{m}$), with $Q_{m}=|Q_{i}|$ (according to Eq. (6)) being the macroscopic charge along the membrane on one side. As $\beta\rightarrow+\infty$, $V_{m}=0$, and the $Q_{m}$ also reaches its maximum as $\gamma\rho_{m}$, with $\rho_{m}|_{\beta\rightarrow+\infty}=2V/(1+\gamma^{2})$. The membrane capacitance $C_{m}=\varepsilon_{m}/d_{0}\rightarrow+\infty$ can be achieved by considering a membrane with zero-thickness $d_{0}\rightarrow 0$ but with finite $\varepsilon_{m}$. From an electrical standpoint, the membrane acts as a simple fluid–fluid interface: without potential jump across the membrane, the Debye layers on both sides still stores charges, and the internal normal traction from internal membrane disappears: $\rho_{m}^{2}\gamma^{2}/\beta|_{\beta\rightarrow 0}=0$. In this limit, the Robin condition in Eq. (15) reduces to the continuity of potential $\phi{]\!]}=0$, and the continuity of the displacement field $\varepsilon_{f}\bm{n\cdot E}{]\!]}=0$ at the interface. These two boundary conditions were used in an EK model for the simple fluid interface [21] that is impermeable to ions. It differs from the interfacial condition using ion conservation [36, 29]. Mechanically, the membrane is different from a simple interface, due to the surface-incompressibility resulting from the conservation of lipids. The limit at $g_{m}\rightarrow 0$ is also discussed in Ref. [10], which represents a short-circuited capacitor and the membrane acts as a simple fluid–fluid interface.
@@
In Ref. [10], the growth rate monotonically increases with $g_{m}$ and then saturates, while in the current system, the growth rate first increases then decreases with $\beta$ and saturates to constant. This is due to the combination effect of internal and external effects. When excluding the internal normal traction (blue curve in Fig. 4(a)), the growth rate behaves monotonically. In Fig. 4(b), the external contribution to the growth rate increases with $\beta$ and saturates at a constant value, whereas the internal traction contribution initially rises from zero at $\beta=0$, peaks, and then decreases back to zero as $\beta\to\infty$. Specifically, $\lim_{\beta\to 0}s_{\text{in}}=\lim_{\beta\to\infty}s_{\text{in}}=0$, corresponding to the overlap between the full growth rate (black curve) and the growth rate excluding internal traction (blue curve) at these limits in Fig. 4(a). The internal traction contribution reaches its maximum at $\beta=\frac{\gamma}{1+\gamma}$, with $\max s_{\text{in}}=\frac{V^{2}\gamma^{3}}{1+\gamma}$. This peak results in a nonmonotonic behavior of the full growth rate in Fig. 4(a), which is further highlighted by the variation of the maximum growth rate shown in Fig. 4(c).
@@
Figure 5: (a) The effect of $\gamma$ on growth rate $s$ as a function of perturbation wavenumber $k$; (b) maximal growth rate $s_{\mathrm{max}}$ as function of $\gamma$. We consider the system with $V=3$, $\beta=1$, $Ca=0.2$ and $\bar{\Gamma}=0.01$.
@@
Fig. 5(a) shows the effect of the asymmetric conductivity $\gamma$ on the growth rate. Due to the reversibility of the system, i.e. $s(\gamma)=s(1/\gamma)$, we only show results for $\gamma>1$. It is very counterintuitive that the membrane subjected to the fluids with the symmetric conductivity ($\gamma=1$) undergoes the most intense instability. As $\gamma$ increases from $1$, the maximum growth rate decreases as shown in Fig. 5(b), and the window for the unstable wavenumbers get narrower. As $\gamma\rightarrow+\infty$, the system becomes stable. This result is very different from the known literature [10], where the instability is amplified as the conductivity becomes more asymmetric. To understand the cause of this difference, it is essential to first examine the mechanism of the instability and how it diverges from the LDM framework. This will be explored in detail in the remainder of this section.
@@ IV.2.1 Stabilizing electrostatic body forces
The Coulomb body forces in the Stokes equation (18), $\bm{f}=-\rho\nabla\phi$, tend to move the positive charges $\rho$ in the direction of the electric field $\bm{E}$. In the base state, these body forces are balanced by the pressure gradient on both sides of the fluid, as evidenced by Eq. (26). Additionally, the flat interface remains free from normal traction arising from either electrostatic stress or elasticity, as well as membrane tension.
(a)(b)(c)(d)(e)(f)
Figure 6: $z$-component of the electric field $E_{z}^{(0)}$ (a), $E_{z}^{(1)}$ (d) and charge density $\rho^{(0)}$ (b), $\rho^{(1)}$ (e), illustrate the components of the first-order electrical body force $\bm{f}^{(1)}\cdot\hat{e}_{z}=\rho^{(0)}E_{z}^{(1)}+\rho^{(1)}E_{z}^{(0)}$ shown in (c,f). The solid curve denotes the perturbed membrane, while the dashed line represents the flat membrane in the base state. Arrows in (c,f) indicate the direction of the $z$-component of the electrical body force. Notably, the force points downward at the crest and upward at the trough of the membrane perturbation, highlighting the restoring nature of the electrical body force. We consider symmetric conductivity $\gamma=1$, $\beta=1$, $V=1$.
To understand the stabilizing terms in Eq. (37) and Eq. (41), we illustrate the first order body force $\bm{f}^{(1)}=-(\rho^{(0)}\nabla\phi^{(1)}+\rho^{(1)}\nabla\phi^{(0)})$ in Fig. 6(c,f), the $z$-component of the electric field $E_{z}^{(0)}$, $E_{z}^{(1)}$ in Fig. 6(a,d), and charge density $\rho^{(0)}$, $\rho^{(1)}$ in Fig. 6(b,e). The electrical body force is calculated with a uni-modal perturbation for wavenumber $k=0.2$, corresponding to a wavelength five times the Debye length, i.e., $x=5$. The perturbation amplitude is set to $h=0.5$. It is expected that the electrical body force acts only on the fluid in the Debye layer, decaying exponentially in the bulk. Fig. 6(c,f) shows that within the Debye layer, the electrical body force points downward at the interface’s crest and upward at the trough. This implies that the electric body force plays a restoring role, consistent with the negative sign in the term Eq. (41) for small $k$.
@@ IV.2.2 Destabilizing external electrostatic surface traction
The external normal traction from the Maxwell stress is
$\partial_{z}\phi^{(0)}\partial_{z}\phi^{(1)}{]\!]}_{z=0}=(\partial_{z}\phi_{1}^{(0)}\partial_{z}\phi_{1}^{(1)}-\partial_{z}\phi_{2}^{(0)}\partial_{z}\phi_{2}^{(1)})|_{z=0}.$ (45)
In LDM framework [10], when the membrane is fully charged, and no current crosses the membrane, resulting in zero Ohmic current in the bulk fluid, the normal traction is zero due to the absence of a base state electric field $\partial_{z}\phi^{(0)}_{i}=-E_{z}^{(0)}=0$. In the current study, the base state electric field is nonzero at the membrane due to the Debye layer $\partial_{z}\phi_{1}^{(0)}|_{z=0}=\partial_{z}\phi_{2}^{(0)}|_{z=0}=\gamma\rho_{m}$. Thus, Eq. (45) is further simplified as
@@
where $\partial_{z}\phi^{(1)}{]\!]}_{z=0}=-h{[\![}\partial_{zz}\phi^{(0)}{]\!]}_{z=0}=h{[\![}\rho^{(0)}{]\!]}_{z=0}$ has been used according to the first order Robin condition in Eq. (32). Consequently, we interpret the external electrostatic surface traction as the net force exerted on the fluid adjacent to the membrane, namely the net charge density $\rho^{(0)}{]\!]}_{z=0}$ moving in the direction of the base state electric field $\partial_{z}\phi_{1}^{(0)}$ (pointing downward), and modulated by the perturbed membrane at (h)(-h). After transforming from Fourier domain to physical domain, we find that the traction has the same formulation but with $h=h(x)$. Then, when $h(x)<0$, the surface traction points downward and vice versa, which renders the membrane unstable to deformations. This implies the destabilizing effect of the external electrostatic normal traction, consistent with the positive sign in the term Eq. (37) for any small $k$.
In addition, we observe that in contrast to the LDM, the presence of a non-zero base state electric field $\bm{E}^{(0)}$ primarily contributes to the instability. In the LDM, a non-zero base state electric field necessitates either bulk current [10, 7] or membrane charging. However, within the EK framework, $\bm{E}^{(0)}$ exists independently of other effects due to the Debye layer.
@@
Figure 7: (a) The uniform charge density in the base state. (b) The first-order perturbed charge density. The total charge density at the membrane, obtained by summing (a) and (b), is shown in (c). Membrane undulations induce a redistribution of charge along the interface, as illustrated in (c). Under a downward electric field, the resulting Maxwell stress exerts upward traction at the crests and downward traction at the troughs of the membrane, indicating its destabilizing effect on membrane shape. The figure shows the example with $\gamma=1$, $\beta=1$, $V=1$.
Intuitively, we can understand such instability through the charge redistribution. Fig. 7 shows the charge distribution in the fluid along the membrane. At the base state (Fig. 7(a)), the charge is uniformly distributed along the membrane. However, when subjected to a small perturbation (Fig. 7(b)), the linear correction causes positive charge accumulation at the troughs on both sides of the membrane, while negative charge accumulates at the crests. As a result (Fig.7(c)), on the upper surface of the membrane, positive charge accumulates denser at the troughs compared to the crests, while on the lower surface of the membrane, negative charge accumulates less at the troughs than at the crests. This accumulation further results in a force which tends to enhance the perturbation. Similar mechanisms of the instability resulting from surface charge redistribution due to membrane undulation are also discussed in Ref. [15]. Different from the present study, the surface charge in Ref. [15] are adsorbed on the interface, and such that compensates the charge distributed in the Debye layer, i.e. each lipid layer is electroneutral.
Next, we aim to understand why the instability is reduced as the conductivity becomes more asymmetric. We consider the scenario where $\gamma\rightarrow 0$, i.e. $R=\frac{n_{1}^{\infty}}{n_{2}^{\infty}}\rightarrow\infty$. In this limit, $\rho^{(0)}_{1}|_{z=0}=\rho_{m}=\frac{2\beta V}{\gamma+(1+\gamma^{2})\beta}\rightarrow 2V$, $\rho^{(0)}_{2}|_{z=0}=-\gamma^{2}\rho_{m}\rightarrow 0$, and the net charge density across the membrane $\rho^{(0)}{]\!]}_{z=0}\rightarrow 2V$. However, the base state field $\partial_{z}\phi_{i}^{(0)}|_{z=0}=\gamma\rho_{m}\rightarrow 0$, resulting in the entire external traction term approaching zero. Therefore, the instability decreases as $\gamma\rightarrow 0$. Similar results can be derived as $\gamma\rightarrow\infty$, where $\rho^{(0)}{]\!]}_{z=0}\rightarrow 2V$ and the base state field $\partial_{z}\phi_{i}|_{z=0}=\gamma\rho_{m}\rightarrow 0$. The result of these two limits is illustrated in Fig. 8, where the destabilizing external surface traction attains its maximum when subjected to the symmetric conductivity.
(a)(b)(c)
Figure 8: The effect of the conductivity ratio $\gamma$ on (a) base state electric field $E_{z}^{(0)}|_{z=0}$, (b) the base state charge density across the membrane, and (c) the external normal surface traction calculated according to Eq. (46). The example is given with $\beta=10$, $V=1$.
@@ IV.2.3 Destabilizing internal electrostatic surface traction
The electric stress induced by the rearranged charge at the membrane has also been used to explain the internal normal traction in Ref. [7], which establishes the system in the LDM framework, and the surface charge is brought to the interface by the bulk current. In their work, the compressive electrostatic stress has the form $\varepsilon_{m}E_{m}^{2}/2$. For a perfectly flat membrane, the electric stress in the membrane is symmetrically balanced on both sides of the membrane. When subjected to a perturbation, the redistributed charge $\Sigma_{\pm}\sim\pm\varepsilon_{m}E_{m}$ results into an unbalanced stress $(\Sigma_{+}^{2}-\Sigma_{-}^{2})/\varepsilon_{m}$. This analysis shares the same situation as ours, where the base state macroscopic charge is $Q_{i}^{(0)}=\pm E_{z}^{(0)}|_{z=0}=\pm\beta dE_{m}^{(0)}$.
@@
where the jump of the first order internal field $\bigg[\!\!\bigg[\partial_{z}\phi_{m}^{(1)}\bigg]\!\!\bigg]_{z=\pm d/2}\sim E_{m}^{(0)}dk^{2}h$ is nonzero due to the perturbation (see Appendix A for expression). Intuitively, we can understand this jump as the virtual charges inside the membrane and driven by the field $E_{m}^{(0)}$. Like the external traction, we consider the deformation in physical domain, which has the same formulation but with $h=h(x)$. At the membrane where $h(x)<0$, the virtual charges are positive and driven by a negative internal field. Thus, the internal surface traction points downward and vice versa, which renders the membrane unstable to deformations.
As discussed around Eq. (37), the internal normal traction in this work is in agreement with the total traction in Ref. [7], where the external traction vanishes in the limit of large $\varepsilon_{m}$ and small membrane conductivity $\sigma_{m}$. In their work, the internal field $E_{m}$ is related to the external field $E$ though the conservation of current, and thus $E_{m}\sim E\sigma_{i}/\sigma_{m}$, and thus small $\sigma_{m}$ indicates a large internal field. In our work, the internal field $E_{m}^{(0)}\sim E_{z}^{(0)}|_{z=0}/(\beta d)$, and thus $\beta d\ll 1$ indicates a large fields ratio $E_{m}^{(0)}/E_{z}|_{z=0}$. Keeping $d$ finite and letting $\beta\rightarrow 0$, we see the destabilizing effect of the normal traction dominated by the internal traction in Eq. (43). This conclusion is in agreement with Refs. [19, 33], where the “inside” contribution is in general dominant when large voltage drop occurs across the membrane, and in our work $\beta\rightarrow 0$ yields the largest transmembrane potential $V_{m}$ from Eq. (24).
Similar to the external surface traction, as $\gamma\rightarrow 0$ or $\gamma\rightarrow+\infty$, the base state external field $E_{z}^{(0)}|_{z=0}\rightarrow 0$ and thus $E_{m}^{(0)}\rightarrow 0$, and the internal traction term vanishes.